A solution contains 0.120 mol H2O and an unknown number of moles of sodium chloride. The vapor pr... A solution contains 0.120 mol H2O and an unknown number of moles of sodium chloride. The vapor pressure of the solution at 30 ?C is 25.7 torr. The vapor pressure of pure water at this temperature is 31.8 torr. Calculate the number of grams of sodium chloride in the solution. (Hint: Remember that sodium chloride is a strong electrolyte.) Mass: g
Answer #1

P(solution) =I* mole fraction of water * pressure of pure water

25.7 / 31.8 = i*mole fraction of water

Mole fraction of water = 0.8082 moles / 2 = 0.4041

Now , mole fraction of water = moles of water / total moles

Suppose mole of NaCl = x

= 0.4041 = 0.120 / ( 0.120 + x )

x = 0.177moles

Mass of water = mole * molecular mass of water

= 0.177 * 18

= 3.186 grams

If you satisfied with the solution please like it thankxx

If you have any query please comment